Data is taken on 773 randomly selected adults who are at least 20 years old. The data is broken down according to sex and age. Suppose one person is randomly selected. Find the following empirical probabilities. Round your answers to three decimal places. (a) Find the probability that they are male. (b) Find the probability that they are between 30 and 39 years old. (c) Find the probability that they are between 30 and 39 or they are older than 40. (d) Find the probability that they are female and older than 40. (e) Find the probability that they are female or older than 40. (f) Find the probability that they are female, given that they are older than 40

Male Female Totals
Between 20 & 29 83 85 168
Between 30 & 39 109 117 226
Older than 40 195 184 379
Totals 387 386 773
Answer #1

a) P(Male) = 387/773 = 0.501

b) P(30 - 39) = 226/773 = 0.292

c) P(30 - 39 or older than 40) = (226 + 379)/773 = 0.783

d) P(Female and 40+) = 184/773 = 0.238

e) P(Female or 40+) = (386 + 195)/773 = 0.752

f) P(Female | 40+) = 184/379 = 0.485

All answers are rounded to 3 decimal places.

I hope these will help. Thank you.